OpenStudy (anonymous):

Given the function g(x) = 4x − 5, find g(−3). −17 7 −7 2 please help meee

Parth (parthkohli):

When a number is inserted in the function \(f\), you get 4 times the number minus 5. Similarly, you have to find four times -3 minues 5.

Parth (parthkohli):

4(-3) - 5 minus*

Parth (parthkohli):

Can you tell me what 4 * -3 is?

OpenStudy (anonymous):

so the answer would be....?

Parth (parthkohli):

4(-3) - 5 solve

OpenStudy (anonymous):

-12

Parth (parthkohli):

yes, now what is -12 - 5?

OpenStudy (anonymous):

-17

Parth (parthkohli):

Yeah, that's it

OpenStudy (anonymous):

oh thanks! :)

Parth (parthkohli):

You're welcome

Parth (parthkohli):

Oh wait it was g not f :P

Parth (parthkohli):

Well anyway the answer is correct :p

hero (hero):

x = -3, therefore: g(x) = 4x − 5 g(-3) = 4(-3)-5 g(-3) = -12 - 5 g(-3) = -17

Parth (parthkohli):

Wow